OpenStudy (superkay):

At a bargain store, Tanya bought 3 items that each cost the same amount.. Tony bought 4 items that cost the same amount , but each was $2.25 less then the items Tonya bought. Both tNya and tony paid the same amount of money. What was the individual cost of each persons items? (A) Write an equation. Let x represent the cost of one of Tanya's items. (B) Solve the equation. (C) Check your solution (D) State the solution

OpenStudy (studygurl14):

3x = 4(x - 2.25)

OpenStudy (studygurl14):

To solve it, your first step would be to distribute the 4

Directrix (directrix):

Tanya: x is the amount per item. 3 items cost: 3*x Tony: y is the amount per item 4 items cost 4*y Tanya and Tony spend the same amount of money for purchased items. That means: 3*x = 4*y But each of Tony's items is $2.25 less than each of Tanya's items. y = x - 2.25 Go back to: 3*x = 4*y 3*x = 4*(x - 2.25) substituting x - 2.25 for y 3x = 4x - 9 Solve for x to get the cost per item for Tanya. @superkay Post what you get. Then, we can finish the problem, okay?

OpenStudy (superkay):

9?

OpenStudy (superkay):

@directrix

OpenStudy (studygurl14):

Yes, 9. Good job

OpenStudy (studygurl14):

So if Tanya's cost $9, how much did Tony's cost?

Directrix (directrix):

Tonys Items: y = 9 - 2.25 y = ? @superkay

OpenStudy (superkay):

6.75

OpenStudy (studygurl14):

right

Directrix (directrix):

Correct. (C) Check your solution Show that this is true: 3*x = 4*y 3* 9 = 4* 6.75 @superkay Show that this is a true statement

Directrix (directrix):

What is 3*9 = What is 4*6.75 =

OpenStudy (superkay):

27

OpenStudy (superkay):

@directrix

OpenStudy (studygurl14):

yes 3*9 = 27

OpenStudy (studygurl14):

and so does 4*6.75 So, you know you're answers are correct